OpenStudy (anonymous):

what is the equation in standard form of a parabola that contains the following points (-2, -20), (0 ,-4),(4,-20)

OpenStudy (anonymous):

Given : What is the equation, in standard form, of a parabola that contains the following points: (–2, 18), (0, 2), (4, 42) ax^2 + bx + c = y - standard form, of a parabola equation Solve for a b and c : Step 1 : Solve for c { this is a give me. you will see why in the below } ax^2 + bx + c = y a(0^2) + b(0) + c = 2 - plugging or replacing x with 0 and y with 2 since anything times zero is 0 so 0 + 0 + c = 2 - the above reasons c = 2 - Addition see I told you this was a give me solve for a ax^2 + bx + 2 = y - Plugging c with 2 a(-2^2 ) - 2b + 2 = 18 - plugging x and y with -2 and 18 a(4^2 ) + 4b + 2 = 42 - plugging x and y with 4 and 42 4a - 2b + 2 = 18 16a + 4b + 2 = 42 - Solving the squares of - 2 and 4 Need a common value to remove the b by multiplication and solve for a though the addition of terms 4 is good so I will use it (2) 4a - 2b + 2 = 18( 2) 16a + 4b + 2 = 42 - multiplying 2 to one of the equation to remove the b's from it by the addition of terms 8a - 4b + 4 = 36 16a + 4b + 2 = 42 --------------------------- - Multiplication and the addition of terms 24a + 6 = 78 24a + 6 = 78 - -derived equation from the above 24a + 6 - 6 = 78 + - 6 -Adding a -6 to both sides of the equation to remove 6 from one side of it and put - 6 on the other side 24a = 72 - Addition 1/24 * 24a = 72 * 1/24 - timing both sides of the equation with 1/24 to solve for a a = 72/24 - multiplication a = 3 - division Let solve for b 4a - 2b + 2 = 18 - one of the equation from the above 4(3) - 2b + 2 = 18 - plugging a with 3 12- 2b + 2 = 18 - multiplication of 4 and 3 14 - 2b = 18 - Addition of 12 and 2 -14 + 14 - 2b = 18 - 14 - -Adding a -14 to both sides of the equation to remove 14 from one side of it and put - 14 on the other side -2b = 4 - Addition -1/2 * -2 b = 4 * -1/2 timing both sides of the equation with -1/2 to solve for b b = -4/2 - multiplication b = - 2 - Division So a = 3 , b = -2 , c = 2 or 3x^2 + -2x + 2 = y ===> your equation ( or Answer) Proof or check : ax^2 + bx + c = y 3x^2 + - 2 x + 2 = y - plugging a , b, c with 3,-2,2 3(-2^2) - 2 (-2) + 2 = 18 - plugging x and y with -2 and 18 3(0^2) -2( 0) + 2 = 2 - plugging x and y with 0 and 2 3(4^2) - 2 (4) + 2 = 42 - plugging x and y with 4 and 42 3(4) + 4 + 2 = 18 3(0) + 0 + 2 = 2 3(16) - 8 + 2 = 42 - solving the squares of -2,0,4 and multiplication 12 + 4 + 2 = 18 0 + 0 + 2 = 2 48 - 8 + 2 = 42 - multiplication 18 = 18 2 = 2 42 = 42 - Addition It checks and equals

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome no problem